OpenStudy (anonymous):

Assume that x and y are positive numbers. Use the properties of logarithms to write the expression in terms of the logarithms of x and y. ln x sqrt{y}

OpenStudy (anonymous):

\[\ln x \sqrt{y}\]

OpenStudy (solomonzelman):

\[\ln(b)=Log_ex\] (in ln(b) the base is ten that's what it is when unspecified just like\[\sqrt{b}~~~is~~~\sqrt[2]{b}\] see?)